You need to accumulate $10,000. To do so, you plan to make deposits of $1,950 per year - with the first payment being made a year from today - into a bank account that pays 8.05% annual interest. Your last deposit will be less than $1,950 if less is needed to round out to $10,000. How many years will it take you to reach your $10,000 goal? How large will the last deposit be?   
Answer #1

Formula for FV of annuity can be used to compute number of periods as:

FV = P x [(1+r) n – 1/r]

P = Periodic cash flow = $ 7,000

r = Rate per period = 8.05 % or 0.0805 p.a.

n = Numbers of periods

$ 10,000 = $ 1,950 x [(1+0.0805) n – 1/0.0805]

$ 10,000 x 0.0805 = $ 1,950 x [(1.0805) n – 1]

$ 805 / $ 1,950 = (1.0805) n – 1)

0.412820513 = (1.0805) n – 1

0.412820513 + 1 = (1.0805) n

(1.0805) n = 1.412820513

Taking logarithm of both sides we get,

n x log 1.0805 = log 1.412820513

n x 0.03362477121 = 0.15008699188

n = 0.15008699188/0.03362477121 = 4.463584033 or 5 periods

Five years need to deposit to accumulate $ 10,000

But the last deposit will be less than $ 1,950.

Let’s compute the last deposit.

Year

Cash Flow (C)

FV Factor computation

FV Factor @ 8.05% (F)

FV

(=C x F)

1

$1,950

(1+0.0805)^4

1.3630101341401

$2,657.869762

2

$1,950

(1+0.0805)^3

1.2614624101250

$2,459.851700

3

$1,950

(1+0.0805)^2

1.1674802500000

$2,276.586488

4

$1,950

(1+0.0805)^1

1.0805000000000

$2,106.975000

5

A

(1+0.0805)^0

1

A

Total PV

$10,000

               

Let the last deposit be A.

Total FV = $2,657.869762 + $2,459.851700 + $2,276.586488 + $2,106.975000 + A

$ 10,000 = $ 9,501.2829488 + A

A = $ 10,000 - $ 9,501.2829488 = $ 498.7170512 or $ 498.72

The last deposit will be of $ 498.72